OpenStudy (anonymous):

raindrops fall 1520m from a coud to the ground. If they were not slowed by air resistance, how fast would the drops be moving when they struck the ground ?

OpenStudy (anonymous):

use v^2=u^2+2as

OpenStudy (anonymous):

where v=final velocity , u=initial velocity , s=height of the raindrop and a =acceleration due to gravity

OpenStudy (anonymous):

\(mgh= \frac 12 mv^2 \implies v=\sqrt{2gh}\)